Skip to Search

 
diagonal spacer

Systems Staff Meeting Notes
March 21, 2008

Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Chris Nicolich, Amy Polk, Mark Sullivan, Stephen Williams
Announcements
  1. Support Services meeting - discussion of the Provost resignation. NOTIS will be turned off in June.
Review Minutes of 3/14/08
  1. Correction: Mario Kart release date is April 27.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. MSL replacement laptop - received, and it works. Should be good to go - they will send us the new ones and them we'll send the old ones back.
  2. Print daemon workstations - to be replaced. Works fine on XP. There are five machines to replace.
Liaisons
  1. Training - desktop support. Not scheduled yet, but it will happen. The whole training is available online. The pre-requisite for the course will be to have attended the other liaison trainings.
  2. Systems training - yesterday. Group C is next.
  3. Achilles - risk assessment training - mandatory for some.
Networking
  1. McAfee - will roll it out next Tuesday.
  2. SAN - user folder migration. We have the capacity to do it now. Will present it to each department. 5G will be allocated for each user folder. User folder data takes up more space than departmental folders. There can be no shared access to user folders. Departmental or project folders will need to be used for sharing. Folders will not be browse-able for security reasons. Will is going to announce the user folder migration at the next Middle Managers meeting. We will set up meetings with each department and present a timeframe in which to move folders. List of departments to come from HR.
  3. CJK - Logan didn't get a chance to meet with David Hickey yet, but he'll try to reach him today and get going on it on Monday.
  4. SAN relocation - no word from Directors.
  5. Central Exchange meeting cancelled at last minute.
  6. Public printing/campus AD/Bridges - no word. Need to get approval from Shelley, then: Judy - to Bridges.
  7. Space for VM backups - have not found the space yet. Once we have more space on NT2R - that should help.
  8. Backup Exec upgraded to new version.
  9. Slip stream - only a couple more critical updates since the last burn - mostly Office updates.
  10. BBS - fighting a battle with porn spam links. Logan turned off anonymous posting on the Minutes topic. Hopefully that will put an end to it.
Software Projects Status
  1. The blog server is coming along, slowly. Logan broke the blog homepage. He needs to re-upload and then migrate the blog from the test server URL.
  2. Grover - ongoing.
  3. Tracking database - no major changes.
  4. Greenstone - Mark is making changes to the toolkit. We send items to external partners and they have to have a license code to FTP to us. It is a bit more secure. Hope to have this finished by Monday.
  5. Circulation Analysis - nothing.
  6. LTMS - tape conversion project - location now designated as VTS. We generated a list of reels that matches the data already migrated. There are over 1200 that can be deleted. Next step will be disposal of old media.
Systems Security Management
  1. Meeting held.

 

University of Florida Home Page